# KACQ-DCNN: Uncertainty-Aware Interpretable Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network for Heart Disease Detection

### v3

@@ Abstract
Heart failure remains a critical global health issue, contributing significantly to cardiovascular disease and accounting for millions of annual deaths. The need for innovative diagnostic strategies is pressing, as classical machine learning models face challenges such as handling complex, high-dimensional data, class imbalances, poor categorical feature representations, limited performance on small datasets, and the absence of uncertainty quantification. Moreover, the interpretability of these models is often hindered by their ‘black box’ nature, complicating clinical trust and decision-making. While quantum machine learning shows potential, existing hybrid models have yet to fully capitalize on quantum advantages. To address these gaps, we propose Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network (KACQ-DCNN), a novel hybrid dual-channel neural network that integrates Kolmogorov–Arnold Networks (KANs) in place of traditional multilayer perceptions, enabling univariate learnable activation functions on edges. As an early adopter of KAN components, we observed that the approach significantly improved the ability of the network to approximate continuous functions with reduced complexity and improved generalizability. Our comprehensive evaluation demonstrates that the KACQ-DCNN 4-qubit 1-layered model outperforms 37 benchmark models, including 16 classical machine learning models, 12 quantum neural networks, six hybrid models, and three variants of the KACQ-DCNN architecture. It achieved an accuracy of 92.03%, along with a macro-average precision, recall, and F1 score of 92.00%, representing significant improvements across all metrics. Moreover, KACQ-DCNN achieved a ROC-AUC score of 94.77%, supported by two-tailed paired t-tests against nine top-performing models, with a significance level () of 5% and a Bonferroni correction applied (), using stratified 10-fold cross-validation. Ablation studies confirmed the synergistic effect of integrating classical-quantum components with KAN, improving accuracy by compared to MLP variants. Furthermore, model-agnostic explainability techniques like LIME and SHAP enhanced feature-level interpretability, and conformal prediction techniques provided robust uncertainty quantification. Our findings indicate that KACQ-DCNN advances prediction accuracy for timely interventions and provides a foundation for more interpretable and trustworthy cardiovascular diagnostics.
@@
Kolmogorov–Arnold Networks (KAN), Classical-Quantum Neural Network, Conformal Prediction, Uncertainty Quantification, Explainable AI, Heart Disease PredictionIntroduction
@@
Heart failure (HF) is a severe health condition in which the heart cannot pump sufficient blood to meet the body’s needs, contributing significantly to global morbidity and mortality [1]. The increase in HF cases is largely attributed to modern dietary habits, sedentary lifestyles, and increased work-related stress [2, 3]. According to the World Health Organization (WHO), cardiovascular disease (CVD) accounts for million deaths annually, with coronary artery disease (CAD) being the most common [3]. CAD is caused by blockages in the coronary arteries that impede blood flow, leading to severe complications such as heart attack and vascular rupture. This condition affects of the global population and consumes a similar percentage of healthcare budgets [2], creating a substantial financial burden while reducing individual productivity. This underscores the critical need for preventive and diagnostic strategies to reduce CAD-related mortality.
@@
Traditional heart disease diagnostics rely on assessing risk factors such as age, sex, family history, lifestyle choices, and symptom analysis [4]. However, this manual approach often fails to capture the hidden parameters and is computationally expensive [5, 2]. Although angiography remains the gold standard for diagnosing CAD [2, 4], its high cost and potential side effects limit its widespread use. Additionally, image-based diagnostic methods are impractical for large-scale screening because of their high cost. In response, researchers are increasingly focusing on developing noninvasive, cost-effective, and rapid automated diagnostic systems for early CAD detection via machine learning (ML) algorithms. Recent advances in artificial intelligence (AI) and ML have significantly improved diagnostic accuracy and accessibility by leveraging large datasets to uncover nonlinear and hierarchical patterns that traditional clinical methods often overlook. These technologies offer a more effective means of analyzing complex medical data and addressing the uncertainty and lack of interpretability that can lead to misdiagnoses or missed cases in conventional approaches.
@@
Current HF prediction models utilizing classical ML (CML) [1, 6, 7, 8, 9, 10, 11], deep learning (DL) [12, 4, 13, 14], hybrid [15, 16, 17], and quantum ML (QML) [2, 18, 19, 20, 21, 22, 23, 24] techniques face several significant research gaps that limit their effectiveness, particularly when applied to short, imbalanced, and tabular datasets characterized by a mix of categorical and numerical features [25]. The integration of classical–quantum ML (CQML) techniques into heart disease prediction remains largely unexplored despite their potential to increase predictive performance through improved optimization and demonstrate computational capabilities even on short and imbalanced datasets [25, 26, 27]. Similarly, while the latest Kolmogorov–Arnold networks (KANs) [28, 29, 30, 31] offer theoretical advantages over MLP-based models and CNNs—such as faster convergence using fewer parameters, learnable univariate functions on edges, optimization capability of activation functions, and smoother approximations with B-splines—they have yet to be fully implemented on practical datasets, limiting their real-world applications such as heart failure detection and investigating its interdisciplinary combinations with fields like QML.
@@
Many ML-DL models operate as “black boxes,” making it challenging for healthcare professionals to understand the rationale behind predictions. This lack of transparency can hinder trust in these systems, especially when critical clinical decisions are at stake [32, 33]. Traditional models often fail to quantify the uncertainty in predictions, which is crucial for clinical decision-making. The integration of uncertainty-aware methods within classical–quantum frameworks could enhance the reliability of predictions; however, this has not been adequately explored. Heart disease datasets often suffer from class imbalance, where instances of heart disease are significantly fewer than those of non-disease cases. Traditional models may not adequately address this issue, leading to biased predictions that favor the majority class[34, 33]. Additionally, these models frequently struggle with overfitting, especially when trained on small datasets, leading to poor generalizability to unseen data[2, 35]. Furthermore, handling categorical features can be inadequate because traditional encoding methods may not effectively capture the underlying relationships. Traditional statistical methods often assume linear relationships between variables, which may not reflect the complex, nonlinear interactions present in cardiovascular health data[32, 35]. Moreover, the significant computational demands and extended training durations of DL models can hinder efforts to enhance the accuracy of heart sound classification [1].
@@
We present the “Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network (KACQ-DCNN),” a hybrid model that combines the unique strengths of the KAN, CQML, explainable AI (XAI) methods such as SHAP and LIME, and uncertainty quantification through conformal prediction techniques. KACQ-DCNN addresses multiple challenges in ML by replacing MLP layers with KAN layers, particularly in managing short, imbalanced, high-categorical, and high-dimensional datasets, offering a novel approach to solving real-world problems such as heart disease prediction. At the core of this innovation is the KAN-enhanced CQML model, which uses learnable activation functions on the edges, allowing it to approximate continuous functions with fewer parameters and reduced complexity. Its spline-based edge functions offer smoother and interpretable mappings, enabling faster convergence and improved generalization [29]. KAN excels in handling high-dimensional, nonlinearly separable data and provides the optimization facility of the activation function, unlike traditional MLPs and CNNs [30, 31]. Furthermore, KACQ-DCNN leverages the exponential computational power of quantum techniques, allowing for the faster exploration of vast solution spaces and the ability to capture detailed correlations and patterns. The quantum components significantly boost scalability, making the model adept at handling small and large datasets while retaining computational efficiency [25]. We selected BiLSTMKANnet as channel-1 and QDenseKANnet as channel-2 after thorough benchmarking, as both models demonstrated superior performance compared to all other evaluated options. Our choice of a dual-channel neural network (DCNN) was driven by its ability to process diverse types of input data in parallel, leveraging distinct feature extraction capabilities for each channel. This approach allows us to combine the individual predictions from each channel to produce final probabilities, effectively enhancing overall performance without compromising efficiency. This architecture enabled us to effectively capture both local and global contexts, which previously worked well with sentiment analysis tasks [36], where understanding nuanced interactions is crucial. Furthermore, it minimizes the risk of overfitting and offers robustness against noise and variability in real-world applications, allowing the model to adapt to dynamic data streams while maintaining accuracy [37, 38]. In addition, incorporating XAI tools such as SHAP and LIME enhances the model’s transparency, offering interpretable insights into predictions, which is especially valuable in fields such as healthcare. To ensure reliability, KACQ-DCNN integrates uncertainty quantification through conformal prediction, provides robust confidence intervals, and addresses ambiguous or high-risk classification scenarios.
@@
Related literature
@@
Various supervised ML algorithms have been utilized to diagnose heart disease. Recent studies have conducted detailed comparative analyses of the effectiveness of these prediction models for heart disease across widely used datasets. Kamala et al. [6] explored the performance of models such as support vector machines (SVM), Random Forest (RF), and K-nearest neighbors (KNN) on heart disease datasets. Among these, RF consistently outperformed the others across several metrics, including sensitivity, accuracy, and specificity. Similarly, Das and Sinha [7] conducted a comprehensive analysis of multiple algorithms, including SVM, Naive Bayes (NB), decision trees (DT), and logistic regression (LR). Their results indicated that SVM and a voting classifier achieved the highest classification accuracy, suggesting that ensemble methods could significantly benefit early diagnosis. Shrivastava et al. [14] developed a hybrid model integrating convolutional neural networks (CNN) and bidirectional long short-term memory (BiLSTM) networks to address issues such as missing data and data imbalance. Their model achieved a 96.66% accuracy on the Cleveland UCI dataset, illustrating the power of combining DL techniques with ML models to enhance performance. Arooj et al. [39] demonstrated the utility of a deep CNN (DCNN) in heart disease prediction, achieving high accuracy, recall, and F1 scores, with a validation accuracy of 91.7%. These results highlight the ability of CNNs to outperform traditional ML models when large datasets are available. Nannapaneni et al. [13] explored a hybrid model that combined CNNs, LSTM networks, and gated recurrent units (GRU), achieving 93.3% accuracy, further demonstrating the potential of DL for heart disease detection. However, the complexity of these models can lead to higher computational costs and longer training times, presenting challenges for their clinical deployment. Sahoo et al. [40] evaluated a range of ML algorithms, including LR, KNN, SVM, NB, DT, RF, and XGBoost. Their study found that RF achieved the highest accuracy of 90.16%. Despite this, the study highlighted that no single model performs best across all datasets, and the choice of the optimal model often depends on the specific characteristics of the dataset, such as the size, class balance, and feature distribution. Rao et al. [41] emphasized that the dataset context is crucial when selecting a model because different models excel under various conditions. This underscores the need to consider model selection criteria carefully, including the trade-off between accuracy and interpretability. Singh and Kumar [11] applied a KNN classifier (k=3) to the Cleveland dataset and achieved 87.00% accuracy, whereas Kavita et al. [15] employed a hybrid decision tree-RF model, improving the accuracy to 88.00%. Mohan et al. [17] introduced the HRFLM model, which balances recall and precision and attains an accuracy of 88.40%. Sharma et al. [10] also achieved a remarkable recall (99.00%) via RF, reinforcing the consistent performance of this algorithm. Saboor et al. [9] reported a 96.72% accuracy with an SVM classifier, further demonstrating the robustness of CMLs on the Cleveland dataset.

### v4

@@ Abstract
Background: Heart failure remains a critical global health issue, contributing significantly to cardiovascular disease burden and accounting for approximately 17.8 million annual deaths worldwide. Traditional diagnostic approaches face substantial limitations in early detection and intervention planning.
Problem: Classical machine learning models struggle with complex, high-dimensional data, class imbalances, poor categorical feature representations, and lack interpretability due to their ‘black box’ nature. While quantum machine learning shows potential, existing hybrid models have yet to capitalize on quantum advantages for cardiovascular diagnostics fully.
Solution: We propose the Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network (KACQ-DCNN), a novel hybrid architecture that integrates KAN components with quantum circuits, enabling univariate learnable activation functions that significantly improve function approximation with reduced complexity. Our comprehensive evaluation demonstrates that the KACQ-DCNN 4-qubit 1-layered model outperforms 37 benchmark models with 92.03% accuracy and 94.77% ROC-AUC score. Ablation studies confirm a synergistic effect of classical-quantum components with KAN, improving accuracy by approximately 2% compared to MLP variants.
Benefits: KACQ-DCNN improves heart disease detection accuracy while providing interpretable insights through model-agnostic explainability techniques like Local Interpretable Model-agnostic Explanations (LIME) and Shapley Additive exPlanations (SHAP). Additionally, conformal prediction techniques deliver robust uncertainty quantification, advancing the field toward more reliable, transparent, and clinically trustworthy cardiovascular diagnostic systems that facilitate timely and effective interventions.
@@
[…]
1 Introduction
@@
Heart failure (HF) is a severe health condition in which the heart cannot pump sufficient blood to meet the body’s needs, contributing significantly to global morbidity and mortality (Rahman et al.,, 2024). The increase in HF cases is largely attributed to modern dietary habits, sedentary lifestyles, and increased work-related stress (Babu et al.,, 2024; Srinivasan et al.,, 2023). According to the World Health Organization (WHO), cardiovascular disease (CVD) accounts for million deaths annually, with coronary artery disease (CAD) being the most common (Srinivasan et al.,, 2023). CAD is caused by blockages in the coronary arteries that impede blood flow, leading to severe complications such as heart attack and vascular rupture. This condition affects of the global population and consumes a similar percentage of healthcare budgets (Babu et al.,, 2024), creating a substantial financial burden while reducing individual productivity. This underscores the critical need for preventive and diagnostic strategies to reduce CAD-related mortality.
@@
Traditional heart disease diagnostics rely on assessing risk factors such as age, sex, family history, lifestyle choices, and symptom analysis (Rao et al.,, 2024). However, this manual approach often fails to capture the hidden parameters and is computationally expensive (Li et al.,, 2020; Babu et al.,, 2024). Although angiography remains the gold standard for diagnosing CAD (Babu et al.,, 2024; Rao et al.,, 2024), its high cost and potential side effects limit its widespread use. Additionally, image-based diagnostic methods are impractical for large-scale screening because of their high cost. In response, researchers are increasingly focusing on developing noninvasive, cost-effective, and rapid automated diagnostic systems for early CAD detection via machine learning (ML) algorithms. Recent advances in artificial intelligence (AI) and ML have significantly improved diagnostic accuracy and accessibility by leveraging large datasets to uncover nonlinear and hierarchical patterns that traditional clinical methods often overlook. These technologies offer a more effective means of analyzing complex medical data and addressing the uncertainty and lack of interpretability that can lead to misdiagnoses or missed cases in conventional approaches.
@@
Current HF prediction models utilizing classical ML (CML) (Rahman et al.,, 2024; Praveena Rachel Kamala et al.,, 2023; Das and Sinha,, 2023; Rani et al.,, 2021; Saboor et al.,, 2022; Sharma et al.,, 2020; Singh and Kumar,, 2020), deep learning (DL) (DeGroat et al.,, 2024; Rao et al.,, 2024; Nannapaneni et al.,, 2023; Shrivastava et al.,, 2023), hybrid (Kavitha et al.,, 2021; Mir et al.,, 2024; Mohan et al.,, 2019), and quantum ML (QML) (Babu et al.,, 2024; Dunjko et al.,, 2016; Wan et al.,, 2017; Rebentrost et al.,, 2014; Schetakis et al.,, 2022; Yoshioka et al.,, 2024; Kumar et al.,, 2021; Kavitha and Kaulgud,, 2023) techniques face several significant research gaps that limit their effectiveness, particularly when applied to short, imbalanced, and tabular datasets characterized by a mix of categorical and numerical features (Jahin et al.,, 2023). The integration of classical–quantum ML (CQML) techniques into heart disease prediction remains largely unexplored despite their potential to increase predictive performance through improved optimization and demonstrate computational capabilities even on short and imbalanced datasets (Jahin et al.,, 2023; Abbas et al.,, 2021; Beer et al.,, 2020). Similarly, while the latest Kolmogorov–Arnold networks (KANs) (Yu et al.,, 2024; Liu et al.,, 2024; Koenig et al.,, 2024; Jamali et al.,, 2024) offer theoretical advantages over MLP-based models and CNNs—such as faster convergence using fewer parameters, learnable univariate functions on edges, optimization capability of activation functions, and smoother approximations with B-splines—they have yet to be fully implemented on practical datasets, limiting their real-world applications such as heart failure detection and investigating its interdisciplinary combinations with fields like QML.
@@
Many ML-DL models operate as “black boxes,” making it challenging for healthcare professionals to understand the rationale behind predictions. This lack of transparency can hinder trust in these systems, especially when critical clinical decisions are at stake (Hossain et al.,, 2024; Baghdadi et al.,, 2023). Traditional models often fail to quantify the uncertainty in predictions, which is crucial for clinical decision-making. The integration of uncertainty-aware methods within classical–quantum frameworks could enhance the reliability of predictions; however, this has not been adequately explored. Heart disease datasets often suffer from class imbalance, where instances of heart disease are significantly fewer than those of non-disease cases. Traditional models may not adequately address this issue, leading to biased predictions that favor the majority class (Al-Alshaikh et al.,, 2024; Baghdadi et al.,, 2023). Additionally, these models frequently struggle with overfitting, especially when trained on small datasets, leading to poor generalizability to unseen data (Babu et al.,, 2024; Subramani et al.,, 2023). Furthermore, handling categorical features can be inadequate because traditional encoding methods may not effectively capture the underlying relationships. Traditional statistical methods often assume linear relationships between variables, which may not reflect the complex, nonlinear interactions present in cardiovascular health data (Hossain et al.,, 2024; Subramani et al.,, 2023). Moreover, the significant computational demands and extended training durations of DL models can hinder efforts to enhance the accuracy of heart sound classification (Rahman et al.,, 2024).
@@
We present the “Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network (KACQ-DCNN),” a hybrid model that combines the unique strengths of the KAN, CQML, explainable AI (XAI) methods such as Shapley Additive exPlanations (SHAP) and Local Interpretable Model-agnostic Explanations (LIME), and uncertainty quantification through conformal prediction techniques. KACQ-DCNN addresses multiple challenges in ML by replacing MLP layers with KAN layers, particularly in managing short, imbalanced, high-categorical, and high-dimensional datasets, offering a novel approach to solving real-world problems such as heart disease prediction. At the core of this innovation is the KAN-enhanced CQML model, which uses learnable activation functions on the edges, allowing it to approximate continuous functions with fewer parameters and reduced complexity. Its spline-based edge functions offer smoother and interpretable mappings, enabling faster convergence and improved generalization (Liu et al.,, 2024). KAN excels in handling high-dimensional, nonlinearly separable data and provides the optimization facility of the activation function, unlike traditional MLPs and CNNs (Koenig et al.,, 2024; Jamali et al.,, 2024). Furthermore, KACQ-DCNN leverages the exponential computational power of quantum techniques, allowing for the faster exploration of vast solution spaces and the ability to capture detailed correlations and patterns. The quantum components significantly boost scalability, making the model adept at handling small and large datasets while retaining computational efficiency (Jahin et al.,, 2023). We selected BiLSTMKANnet as channel-1 and QDenseKANnet as channel-2 after thorough benchmarking, as both models demonstrated superior performance compared to all other evaluated options. Our choice of a dual-channel neural network (DCNN) was driven by its ability to process diverse types of input data in parallel, leveraging distinct feature extraction capabilities for each channel. This approach allows us to combine the individual predictions from each channel to produce final probabilities, effectively enhancing overall performance without compromising efficiency. This architecture enabled us to effectively capture both local and global contexts, which previously worked well with sentiment analysis tasks (Zhang et al.,, 2023), where understanding nuanced interactions is crucial. Furthermore, it minimizes the risk of overfitting and offers robustness against noise and variability in real-world applications, allowing the model to adapt to dynamic data streams while maintaining accuracy (Ye et al.,, 2024; Zheng et al.,, 2024). In addition, incorporating XAI tools such as SHAP and LIME enhances the model’s transparency, offering interpretable insights into predictions, which is especially valuable in fields such as healthcare (Dhar et al.,, 2023). To ensure reliability, KACQ-DCNN integrates uncertainty quantification through conformal prediction, provides robust confidence intervals, and addresses ambiguous or high-risk classification scenarios.
@@
2 Related literature
2.1 Supervised ML and DL Models for Heart Disease Prediction
@@
Various supervised ML algorithms have been utilized to diagnose heart disease. Recent studies have conducted detailed comparative analyses of the effectiveness of these prediction models for heart disease across widely used datasets. Kamala et al. (Praveena Rachel Kamala et al.,, 2023) explored the performance of models such as support vector machines (SVM), Random Forest (RF), and K-nearest neighbors (KNN) on heart disease datasets. Among these, RF consistently outperformed the others across several metrics, including sensitivity, accuracy, and specificity. Similarly, Das and Sinha (Das and Sinha,, 2023) conducted a comprehensive analysis of multiple algorithms, including SVM, Naive Bayes (NB), decision trees (DT), and logistic regression (LR). Their results indicated that SVM and a voting classifier achieved the highest classification accuracy, suggesting that ensemble methods could significantly benefit early diagnosis. Shrivastava et al. (Shrivastava et al.,, 2023) developed a hybrid model integrating convolutional neural networks (CNN) and bidirectional long short-term memory (BiLSTM) networks to address issues such as missing data and data imbalance. Their model achieved a 96.66% accuracy on the Cleveland UCI dataset, illustrating the power of combining DL techniques with ML models to enhance performance. Arooj et al. (Arooj et al.,, 2022) demonstrated the utility of a deep CNN (DCNN) in heart disease prediction, achieving high accuracy, recall, and F1 scores, with a validation accuracy of 91.7%. These results highlight the ability of CNNs to outperform traditional ML models when large datasets are available. Nannapaneni et al. (Nannapaneni et al.,, 2023) explored a hybrid model that combined CNNs, LSTM networks, and gated recurrent units (GRU), achieving 93.3% accuracy, further demonstrating the potential of DL for heart disease detection. However, the complexity of these models can lead to higher computational costs and longer training times, presenting challenges for their clinical deployment. Sahoo et al. (Kumar Sahoo et al.,, 2022) evaluated a range of ML algorithms, including LR, KNN, SVM, NB, DT, RF, and XGBoost. Their study found that RF achieved the highest accuracy of 90.16%. Despite this, the study highlighted that no single model performs best across all datasets, and the choice of the optimal model often depends on the specific characteristics of the dataset, such as the size, class balance, and feature distribution.
Rao et al. (Rao et al.,, 2021) emphasized that the dataset context is crucial when selecting a model because different models excel under various conditions. This underscores the need to consider model selection criteria carefully, including the trade-off between accuracy and interpretability. Singh and Kumar (Singh and Kumar,, 2020) applied a KNN classifier (k=3) to the Cleveland dataset and achieved 87.00% accuracy, whereas Kavita et al. (Kavitha et al.,, 2021) employed a hybrid decision tree-RF model, improving the accuracy to 88.00%. Mohan et al. (Mohan et al.,, 2019) introduced the HRFLM model, which balances recall and precision and attains an accuracy of 88.40%. Sharma et al. (Sharma et al.,, 2020) also achieved a remarkable recall (99.00%) via RF, reinforcing the consistent performance of this algorithm. Saboor et al. (Saboor et al.,, 2022) reported a 96.72% accuracy with an SVM classifier, further demonstrating the robustness of CMLs on the Cleveland dataset.
Study | Method(s) | Dataset | Accuracy | Highlights |
Mir et al., (2024) | RF + Adaboost | Heart Statlog, UCI | Statlog: 99.48%, UCI: 93.90% | Outperformed previous studies |
Praveena Rachel Kamala et al., (2023) | SVM, RF, KNN | Heart Disease DB | RF: 82.11% | RF performed best in sensitivity and accuracy |
Das and Sinha, (2023) | SVM, NB, DT, LR, Voting | Heart Disease DB | SVM/Voting: Best | Ensemble learning was beneficial |
Shrivastava et al., (2023) | CNN + BiLSTM (Hybrid) | Cleveland | 96.66% | Robust to missing data & imbalance |
Arooj et al., (2022) | DCNN | Heart Disease DB | 91.70% | Strong CNN performance |
Nannapaneni et al., (2023) | CNN, LSTM, GRU (Hybrid) | Cleveland | 93.30% | High performance from hybrid DL |
Kumar Sahoo et al., (2022) | LR, KNN, SVM, NB, DT, RF, XGBoost | Cleveland | RF: 90.16% | RF was overall best |
Singh and Kumar, (2020) | KNN (k=3) | Cleveland | 87% | Moderate accuracy with simple method |
Kavitha et al., (2021) | Hybrid DT + RF | Cleveland | 88% | Hybrid model improved results |
Mohan et al., (2019) | HRFLM | Cleveland | 88.40% | Balanced precision and recall |
Sharma et al., (2020) | SVM, RF, DT, NB | Cleveland | – | RF had better recall |
Saboor et al., (2022) | SVM | Cleveland | 96.72% | High accuracy via SVM |
Li et al., (2020) | FCMIM + Classifier | Cleveland | 92% | Boosted recall with feature selection |
Doppala et al., (2023) | Genetic Algorithm + RBF | Cleveland | 94% | Achieved high accuracy with optimization |
Rani et al., (2021) | Random Forest | Cleveland | 86.6% | Balanced metrics |
Shah et al., (2020) | KNN (k=7) | Cleveland | 90.78% | Tuned K produced high accuracy |
2.2 QML in Medical Prediction